### v1

@@ Abstract
In this work, we survey recent studies on masked image modeling (MIM), an approach that emerged as a powerful self-supervised learning technique in computer vision. The MIM task involves masking some information, e.g. pixels, patches, or even latent representations, and training a model, usually an autoencoder, to predicting the missing information by using the context available in the visible part of the input. We identify and formalize two categories of approaches on how to implement MIM as a pretext task, one based on reconstruction and one based on contrastive learning. Then, we construct a taxonomy and review the most prominent papers in recent years. We complement the manually constructed taxonomy with a dendrogram obtained by applying a hierarchical clustering algorithm. We further identify relevant clusters via manually inspecting the resulting dendrogram. Our review also includes datasets that are commonly used in MIM research. We aggregate the performance results of various masked image modeling methods on the most popular datasets, to facilitate the comparison of competing methods. Finally, we identify research gaps and propose several interesting directions of future work.
@@ 1 Introduction
Figure 1: A timeline with the most prominent works in Masked Image Modeling.
@@
Data have always played a crucial role in training large deep neural networks. Procuring a curated labeled dataset not only involves a great effort and a laborious process from a team of human annotators, but it also represents a significant expense. As a result, various self-supervised learning strategies have been explored, where the model is pre-trained with a different objective, which does not require human labels. Self-supervised learning can help the model to learn a rich feature representation, and even surpass supervised alternatives [1]. Then, during the fine-tuning phase, the model is further optimized for a specific downstream task. For example, in most image classification tasks, a common approach is to train the network on the ImageNet dataset [2, 3], and then change the classification head for a new task and train the resulting model on a relevant dataset.
@@
Self-supervised learning is a popular method to pre-train a deep learning model. It involves creating a supervised setting without annotating the data, but rather using the inherent structure of the data. Due to its potential, self-supervised pre-training algorithms have been rapidly adopted in vision in recent years. The early works in this direction by [4] and [5] were inspired by the jigsaw puzzle idea, in which they proposed to split the image into patches and estimate the position of each patch. Another pre-training strategy was introduced by [6], where the objective was to have the distance between the initial and last frame in a video smaller than the distance between the initial frame and a random frame from another video. [7] created a segmentation map from the motion of objects in videos, and then applied a segmentation task for pre-taining, using the artificial segmentation map as ground-truth. Other prominent pretext tasks are colorizing a gray-scale image [8] or rotating the input image and estimating the angle it was rotated with [9].
@@
The idea of pre-training a model by masking part of the input and then predicting the masked information gained traction with the introduction of Bidirectional Encoder Representations from Transformers (BERT) [10], which brought significant advancements in the Natural Language Processing domain. The main advantage was that huge amounts of unlabeled and unstructured text data could be used. Notably, this pretext strategy was applied in vision problems earlier, by [11] and [12], where the original input signal was altered (corrupted or obscured) and then reconstructed using an architecture based on convolutional layers. The later work even employed an encoder-decoder model to inpaint the missing regions from the input image. Nevertheless, as presented by [1], there are some differences from the recent Masked Image Modeling (MIM) literature, as the aforementioned papers framed the problem as a denoising task.
@@
In Figure 1, we present a timeline with the most prominent works in masked image modeling. At ICCV 2021, [13] presented a self-supervised learning method that applies a contrastive loss between a masked patch from an image and other regions. At NeurIPS 2021, [14] proposed a teacher-student framework that employs both a reconstruction and a contrastive objective. At CVPR 2022, [1] and [15] introduced a pre-training framework that involves masking a high portion of the input and reconstruct it using autoencoders based on the Vision Transformer (ViT) architecture [16]. Their concurrent studies represent the base for the research that followed. Following [14], [17] formally presented a pre-training method that involves both objectives, but follows the masked autoencoder (MAE) framework. Besides employing MAE for pre-training, [18] demonstrated how to use MAE at inference time, improving performance on downstream tasks. At ICLR 2023, [19] combined MIM with denoising contrastive learning for a better feature learning, while [20] analyzed the teacher-student MIM framework and showed the advantages of updating the teacher’s weights as an exponential moving average of the student’s. [21] is another important stepping stone in integrating both reconstruction and contrastive objectives in the MAE framework. At NeurIPS 2023, [22] presented a method that applies the masking pre-training method on multiple input vision modalities, as well as text.
@@
Within the context of more complex architectures of the latest neural networks, and the large quantities of annotated data they require, pre-training such models has started to become a prerequisite. Masked Image Modeling represents a pretext task that consists of masking some information from the input (either the raw signal or some features obtained from it), and then estimating an output that should be the same as if the input was unaltered, or even predicting the original input. This pre-training strategy has quickly become popular, especially since the mechanism is easily implemented with the well-known transformer architecture, and thus it emerged in many domains and tasks. As a result, it is very difficult and time-consuming to study such a high number of research papers and find the necessary information. Our work aims to diminish this challenge and facilitate further research or industrial endeavors. Firstly, we present a generic framework that all masked image modeling methods follow and identify two different categories of approaches: one involving input reconstruction and one employing a contrastive objective. Furthermore, we have carefully reviewed the most recent papers and extracted the main ideas and contributions from these studies. Furthermore, we manually organize the reviewed studies into a taxonomy based on multiple criteria, which is complemented by a dendrogram obtained via hierarchical clustering.
@@
Given the increasing prominence of masked image modeling and the corresponding rise in research publications on this topic, several studies have been conducted with objectives similar to ours: to facilitate the literature review process. Among these, the survey on masked modeling by [23] stands out as a notable contribution, with which our paper shares many characteristics, such as the general frameworks employed during pre-training, some criteria of the manual taxonomy, or even some presented papers. Nevertheless, our survey is focused on vision and how MIM is applied in the most recent techniques. The work of [23] surveys a broad range of domains, while including a limited number of papers per domain. Furthermore, the more focused scope of our survey on the image domain allows us to review a greater number of vision papers for a more thorough study. Thus, we consider our work to be more comprehensive, particularly in the computer vision domain. In contrast to [23], we organize the papers via both manual and automatic clustering, providing complementary ways to categorize the surveyed papers.
@@ 2 Generic Framework
Masked Image Modeling is an unsupervised technique that is usually applied during the pre-training phase. It involves masking some information, either from the input or from the latent representation, and then estimating the original data, as if the data would not have been concealed. Although many masked image modeling techniques have been proposed, the research has been focused on two main schemes, either reconstructing the masked signal, or comparing two latent representations, one for the unaltered input signal and one for the masked input. On a few occasions, different approaches are explored, but they are built on similar grounds. Therefore, in the following subsections, we aim to give a general formulation of the first two aforementioned schemes.
@@ 2.1 Reconstruction
We formally present the reconstruction-based training strategy in Algorithm 1. The first step of the algorithm splits the input images into non-overlapping patches, resulting in a set $P$ that contains patches of the same size, namely of $h\times w$ pixels. The second step applies the masking operation, which can differ from one method to another. The masking operation indicates which patches should be kept and which should be masked. The indexes of the visible and masked patches are stored in $I_{v}$ and $I_{m}$, respectively. In the next steps, the masked patches are usually dropped, and only the visible patches are processed by a visual projection layer ($V\!P_{\varphi})$, followed by an encoder ($E_{\theta}$) which extracts a latent representation. Before the decoding step, the previously masked patches are replaced by a learnable representation ($M$), which resides in the latent space of the encoder. These transformations correspond to steps 3-7 in Algorithm 1.
@@ 2.2 Contrastive
Input: $X$ – the input image; $h,w$ – the patch dimensions; $\alpha$ – the proportion of input masking; split – the function that splits an image into a number of patches; mask – the function which chooses what patches should be masked; augment – the image augmentation function; $M$ – the learnable embedding of the masked patches; $\eta$ – the learning rate.
@@
12:$\mathcal{L}\left(\varphi,\!\theta,\!M\right)\!\leftarrow\!-\!\sum_{i=1}^{n}\!\log{\frac{exp(\langle\hat{H}[i],\tilde{H}[i]\rangle/\tau)}{\sum_{j=1}^{n}exp(\langle\hat{H}[j],\tilde{H}[i]\rangle/\tau)}}$
@@
The generic contrastive-based MIM framework is formalized in Algorithm 2. The aim of this framework is to obtain a similar embedding, irrespective of the applied masking. Steps 1 and 2 of the algorithm generate two versions of the input image by augmenting them at different intensities. The version that undergoes masking is strongly augmented, while the other one is weakly augmented (or can even remain unaltered). In steps 3-4, each image is divided into non-overlapping patches, and in step 5, masking is applied solely to the strongly augmented image. The unmasked image undergoes processing by the projection layer and the encoder (step 7), whereas the masked image has its omitted patches replaced with a learnable vector before being encoded (steps 8-11). Notably, Algorithm 2 processes both input images using the same encoder, which can be regarded as two encoders with shared weights. Gradient propagation is restricted to the processing of the masked image. An alternative approach [20] is to process the unmasked image with an EMA-based encoder in steps 9 and 11. Step 12 of the algorithm computes the contrastive loss, in which the negative patches originate from different positions within the same image. Moreover, in practice, the negatives can also be sourced from other images. Finally, the last steps (13-15) of the algorithm update the weights of the encoder, the projection layer, and the learnable representation $M$. Although the contrastive learning method is technically different from the reconstruction-based scheme, [26] demonstrated that the contrastive approach strongly correlates with the reconstruction-based framework in terms of the learned latent representations.
Figure 4: A multi-level classification of Masked Image Modeling papers into various categories, based on the research directions studied in the respective papers.
@@ 3 Taxonomy and Overview
In Figure 4, we present a manually-generated taxonomy of the most promising MIM papers, organizing them according to their main contributions. In constructing the taxonomy, we consider six main research directions related to: the masking strategy, the type of masked signals, the neural architecture, the objective function, the type of downstream tasks, and theoretical results. Nonetheless, since many papers focused on more than one of the above aspects, we grouped the respective works in distinct categories representing composed contributions. We now continue by presenting the aforementioned papers, divided into the sub-sections according our taxonomy.
@@ 3.1 Masking Strategy
Early masking strategies were based on selecting a high number of random patches and masking them. Recently, many alternative masking policies have been proposed to improve MIM. One suggestion is to guide the masking through different mechanisms, from leveraging image statistics [27] to using an additional simple network [28] or even a self-attention module [29, 30].
Figure 5: SimMIM pipeline as presented by [15].
@@
Figure 6: MAE pipeline as presented by [1].
@@
[1] presented the masked autoencoder (MAE) framework, where the main contribution is to exclude the masked tokens from the encoder’s input and to use a very high masking ratio (75%). These changes imply a more efficient framework compared with other works, such as SimMIM [15]. However, as a negative effect, having an encoder that operates only on visible tokens makes the framework incompatible by default with the Hierarchical Vision Transformer [31] architecture, which usually performs better than a standard ViT. The MAE pipeline is presented in Figure 6.

### v2

@@ Abstract
In this work, we survey recent studies on masked image modeling (MIM), an approach that emerged as a powerful self-supervised learning technique in computer vision. The MIM task involves masking some information, e.g. pixels, patches, or even latent representations, and training a model, usually an autoencoder, to predicting the missing information by using the context available in the visible part of the input. We identify and formalize two categories of approaches on how to implement MIM as a pretext task, one based on reconstruction and one based on contrastive learning. Then, we construct a taxonomy and review the most prominent papers in recent years. We complement the manually constructed taxonomy with a dendrogram obtained by applying a hierarchical clustering algorithm. We further identify relevant clusters via manually inspecting the resulting dendrogram. Our review also includes datasets that are commonly used in MIM research. We aggregate the performance results of various masked image modeling methods on the most popular datasets, to facilitate the comparison of competing methods. Finally, we identify research gaps and propose several interesting directions of future work. We supplement our survey with the following public repository containing organized references: https://github.com/vladhondru25/MIM-Survey.
@@ 1 Introduction
Figure 1: A timeline with the most prominent works in Masked Image Modeling. This timeline illustrates the evolution of MIM methods and the primary research directions in this area. Early works adapted masked language modeling directly to the image domain, as in MST [1] and SimMIM [2]. Subsequent efforts [3, 4] tried to combine the MIM ideas with another successful self-supervised method applied in computer vision, namely contrastive learning. The most recent developments extend MIM to multimodal settings [5, 6] and tackle more complex tasks [7, 8].
@@
Data samples have always played a crucial role in training large deep neural networks. Procuring a curated labeled dataset not only involves a great effort and a laborious process from a team of human annotators, but it also represents a significant expense. As a result, various self-supervised learning strategies have been explored, where the model is pre-trained with a different objective, which does not require human labels. Self-supervised learning can help the model to learn a rich feature representation, and even surpass supervised alternatives [9]. Then, during the fine-tuning phase, the model is further optimized for a specific downstream task. For example, in most image classification tasks, a common approach is to train the network on the ImageNet dataset [10, 11], and then change the classification head for a new task and train the resulting model on a relevant dataset.
@@
Self-supervised learning is a popular method to pre-train a deep learning model. It involves creating a supervised setting without annotating the data, but rather using the inherent structure of the data. Due to its potential, self-supervised pre-training algorithms have been rapidly adopted in computer vision in recent years. The early works in this direction by [12] and [13] were inspired by the jigsaw puzzle idea, in which the authors proposed to split the image into patches and estimate the position of each patch. Another pre-training strategy was introduced by [14], where the objective was to have the distance between the initial and last frame in a video smaller than the distance between the initial frame and a random frame from another video. [15] created a segmentation map from the motion of objects in videos, and then applied a segmentation task for pre-taining, using the artificial segmentation map as ground-truth. Other prominent pretext tasks are colorizing a gray-scale image [16] or rotating the input image and estimating the angle it was rotated with [17].
@@
The idea of pre-training a model by masking a part of the input and then predicting the masked information gained traction with the introduction of Bidirectional Encoder Representations from Transformers (BERT) [18], which brought significant advancements in the Natural Language Processing domain. The main advantage was that huge amounts of unlabeled and unstructured text data could be used. Notably, this pretext strategy was applied in vision problems earlier, by [19] and [20], where the original input signal was altered (corrupted or obscured) and then reconstructed using an architecture based on convolutional layers. The latter work even employed an encoder-decoder model to inpaint the missing regions from the input image. Nevertheless, as presented by [9], there are some differences from the recent Masked Image Modeling (MIM) literature, as the aforementioned papers framed the problem as a denoising task.
@@
In Figure 1, we present a timeline with the most prominent works in masked image modeling. At ICCV 2021, [21] presented a self-supervised learning method that applies a contrastive loss between a masked patch from an image and other regions. At NeurIPS 2021, [1] proposed a teacher-student framework that employs both a reconstruction and a contrastive objective. At CVPR 2022, [9] and [2] introduced a pre-training framework that involves masking a high portion of the input and reconstruct it using autoencoders based on the Vision Transformer (ViT) architecture [22]. Their concurrent studies represent the base for the research that followed. Following [1], [23] formally presented a pre-training method that involves both objectives, but follows the masked autoencoder (MAE) framework. Besides employing MAE for pre-training, [24] demonstrated how to use MAE at inference time, improving performance on downstream tasks. At ICLR 2023, [3] combined MIM with denoising contrastive learning for a better feature learning, while [25] analyzed the teacher-student MIM framework and showed the advantages of updating the teacher’s weights as an exponential moving average of the student’s. [4] is another important stepping stone in integrating both reconstruction and contrastive objectives in the MAE framework. At NeurIPS 2023, [26] presented a method that applies the masking pre-training method on multiple input vision modalities, as well as text. More recent contributions are focused on multimodal settings [5, 6] and more complex tasks [7, 8].
@@
Within the context of more complex architectures of the latest neural networks and the large quantities of annotated data they require, pre-training such models has started to become a prerequisite. Masked Image Modeling represents a pretext task that consists of masking some information from the input (either the raw signal or some features obtained from it), and then estimating an output that should be the same as if the input was unaltered, or even predicting the original input. This pre-training strategy has quickly become popular, especially since the mechanism is easily implemented with the well-known transformer architecture, and thus it emerged in many domains and tasks. As a result, it is very difficult and time-consuming to study such a high number of research papers and find the necessary information. Our work aims to mitigate this challenge and facilitate further research or industrial endeavors. Firstly, we present a generic framework that all masked image modeling methods follow and identify two different categories of approaches: one involving input reconstruction and one employing a contrastive objective. Furthermore, we have carefully reviewed the most recent papers and extracted the main ideas and contributions from these studies. Furthermore, we manually organize the reviewed studies into a taxonomy based on multiple criteria, which is complemented by a dendrogram obtained via hierarchical clustering.
@@
Given the increasing prominence of masked image modeling and the corresponding rise in research publications on this topic, several studies have been conducted with objectives similar to ours: to facilitate the literature review process. Among these, the survey on masked modeling by [27] stands out as a notable contribution, with which our paper shares many characteristics, such as the general frameworks employed during pre-training, some criteria of the manual taxonomy, or even some presented papers. Nevertheless, our survey is focused on vision and how MIM is applied in the most recent techniques. The work of [27] surveys a broad range of domains, while including a limited number of papers per domain. Furthermore, the more focused scope of our survey on the image domain allows us to review a greater number of vision papers for a more thorough study. Thus, we consider our work to be more comprehensive, particularly in the computer vision domain. In contrast to [27], we organize the papers via both manual and automatic clustering, providing complementary ways to categorize the surveyed papers. Other relevant surveys are the ones conducted by [28] and [29], where the authors focus specifically on masked autoencoders. Unlike these related surveys, we address the broader domain of masked image modeling, which is not necessarily coupled with autoencoders. Our survey is therefore more comprehensive, containing more than 100 additional references compared with both surveys on masked autoencoders [28, 29]. For example, we review works that integrate masking into basic operations, e.g. convolutions [30, 31]. As shown by the respective authors, masked convolutions can be integrated into any architecture.
@@ 2 Generic Framework
Masked Image Modeling is an unsupervised technique that is usually applied during the pre-training phase. It involves masking some information, either from the input or from the latent representation, and then estimating the original data, as if the data would not have been concealed. Although many masked image modeling techniques have been proposed, the research has been focused on two main schemes, either reconstructing the masked signal, or comparing two latent representations, one for the unaltered input signal and one for the masked input. On a few occasions, different approaches have been explored, but they are built on similar grounds. Therefore, in the following subsections, we aim to give a general formulation of the first two aforementioned schemes.
@@ 2.1 Reconstruction
We formally present the reconstruction-based training strategy in Algorithm 1. The first step of the algorithm splits the input images into non-overlapping patches, resulting in a set $P$ that contains patches of the same size, namely of $h\times w$ pixels. The second step applies the masking operation, which can differ from one method to another. The masking operation indicates which patches should be kept and which should be masked. The indexes of the visible and masked patches are stored in $I_{v}$ and $I_{m}$, respectively. In the next steps, the masked patches are usually dropped, and only the visible patches are processed by a visual projection layer ($V\!P_{\varphi})$, followed by an encoder ($E_{\theta}$) that extracts a latent representation. Before the decoding step, the previously masked patches are replaced by a learnable representation ($M$), which resides in the latent space of the encoder. These transformations correspond to steps 3-7 in Algorithm 1.
@@ 2.2 Contrastive
Input: $X$ – the input image; $h,w$ – the patch dimensions; $\alpha$ – the proportion of input masking; split – the function that splits an image into a number of patches; mask – the function which chooses what patches should be masked; augment – the image augmentation function; $M$ – the learnable embedding of the masked patches; $d$ – the optimization function; $\eta$ – the learning rate.
@@
12:$\mathcal{L}\left(\varphi,\!\theta,\!M\right)\!\leftarrow\!d(\hat{H},\tilde{H})$
@@
The generic contrastive-based MIM framework is formalized in Algorithm 2. The aim of this framework is to obtain a similar embedding, irrespective of the applied masking. Steps 1 and 2 of the algorithm generate two versions of the input image by augmenting them at different intensities. The version that undergoes masking is strongly augmented, while the other one is weakly augmented (or can even remain unaltered). In steps 3-4, each image is divided into non-overlapping patches, and in step 5, masking is applied solely to the strongly augmented image. The unmasked image undergoes processing by the projection layer and the encoder (step 7), whereas the masked image has its omitted patches replaced with a learnable vector before being encoded (steps 8-11). Notably, Algorithm 2 processes both input images using the same encoder, which can be regarded as two encoders with shared weights. Gradient propagation is restricted to the processing of the masked image. An alternative approach [25] is to process the unmasked image with an EMA-based encoder in steps 9 and 11. Step 12 of the algorithm computes the contrastive loss, in which the negative patches originate from different positions within the same image. The most common loss function used in Step 12 is some variant of the InfoNCE (Noise-Contrastive Estimation) loss:
$\mathcal{L}\left(\varphi,\!\theta,\!M\right)\!\leftarrow\!-\!\sum_{i=1}^{n}\!\log{\frac{\exp(\langle\hat{H}[i],\tilde{H}[i]\rangle/\tau)}{\sum_{j=1}^{n}\exp(\langle\hat{H}[j],\tilde{H}[i]\rangle/\tau)}},$ (1)
where $\tau$ is the temperature factor scaling the pairwise similarity. The goal of InfoNCE is to maximize the similarity between $\hat{H}[i]$ and $\tilde{H}[i]$, while minimizing the similarities $\hat{H}[j]$ and $\tilde{H}[i]$, $\forall j\neq i$. In practice, the negative patches denoted by $\hat{H}[j]$ can also be sourced from other images. Finally, the last steps (13-15) of the algorithm update the weights of the encoder, the projection layer, and the learnable representation $M$. Although the contrastive learning method is technically different from the reconstruction-based scheme, [34] demonstrated that the contrastive approach strongly correlates with the reconstruction-based framework in terms of the learned latent representations.
Figure 4: A multi-level classification of Masked Image Modeling papers into various categories, based on the research directions studied in the respective papers. References are clickable links to papers.
@@ 3 Taxonomy and Overview
In Figure 4, we present a manually-generated taxonomy of the most promising MIM papers, organizing them according to their main contributions. In constructing the taxonomy, we consider six main research directions related to: the masking strategy, the type of masked signals, the neural architecture, the objective function, the type of downstream tasks, and theoretical results. Nonetheless, since many papers focused on more than one of the above aspects, we grouped the respective works into distinct categories representing composed contributions. We now continue by presenting the aforementioned papers, divided into sub-sections according to our taxonomy.
@@ 3.1 Masking Strategy
Figure 5: Types of masking strategies employed in MIM pipelines. The masking strategies can either rely on random patch selection or leverage some additional information to selectively mask the salient patches. Additional information can come from auxiliary neural networks or from features derived through classical computer vision algorithms.
In Figure 5, we categorize the main masking strategies employed in MIM pipelines. A common practice is to mask the content randomly [2, 9, 35, 36, 37, 38, 39]. Nevertheless, several recent works advocate for a more informed masking [40, 41, 42, 43, 44, 45, 46, 47, 48, 49, 50], guided by meaningful information, whether it stems from a neural network [41, 42, 51] or from classical features [40, 43, 44, 52]. By masking the salient patches or semantically relevant regions, these informed approaches enhance the capacity of the model to learn robust representations.
Figure 6: SimMIM pipeline, courtesy of [2] (image licensed under CC BY 4.0). SimMIM is a reconstruction-based method (belonging to the generic pipeline in Figure 2), as it employs an ${L}^{1}$ loss between the masked patches and their counterparts from the original image.
@@
Figure 7: MAE pipeline, courtesy of [9] (image licensed under CC BY 4.0). MAE is a reconstruction-based method (belonging to the generic pipeline in Figure 2), as it employs an ${L}^{2}$ loss between the reconstructed patches and their counterparts from the original image.
@@
[9] presented the masked autoencoder (MAE) framework, where the main contribution is to exclude the masked tokens from the encoder’s input and use a very high masking ratio (75%). These changes imply a more efficient framework compared with other works, such as SimMIM [2]. However, as a negative effect, having an encoder that operates only on visible tokens makes the framework incompatible by default with the Hierarchical Vision Transformer [53] architecture, which usually performs better than a standard ViT. The MAE pipeline is presented in Figure 7. Aiming to learn more robust representations, OmniMAE, introduced by [38], is similar to the standard MAE framework, except that the model is trained with both video and image data. Another extension is MixMAE [39], a pre-training approach that leverages elements of both MAE [9] and SimMIM [2]. This method selects two distinct images from a training dataset and extracts random tokens from each. The tokens are then amalgamated to form a new, hybrid image. This composite image undergoes processing by an encoder. Further, a decoder is employed, aiming to reconstruct the original two images from which the mixed image was derived. During the decoding phase, the input provided to the decoder is unmixed and the patches corresponding to the missing tokens in the hybrid image are filled with masked tokens.